Problem 17 (s.oo pts) nd of a 4-oom long metal bar. A mierophone at the other end picks up two pulses of ed by 1.o ms. What is the bar was first sound, one that travels through the bar, the other through the air. The pulses are sep the speed of sound in the bar? Assume the speed of sound in air is 343 m/s and the pulse fron oll s A) 5174 m/s ) 176 m/s C)2587 m/s D) 352 ms E) Cannot be determined Problem 18: (5.oo pts): A traveling wave has a displacement given by y(x,t) = 2.0 an X cos(2tx-41) where x is measured in cm and t in s. What is the speed of the wave? A) 1.00 cm/s 2.00 cm/s 4.00 cm/s D) 6.00 cm/s E) 0.50 cms Problem 19: (5.00 pts): Light from a sodium lamp (λ = 589 nm) illuminates two narrow slits. The spacing between two consecutive brightfrnges is on a screen iso em behind the slits is 2.o mm. What is the spacing (Gn mm) between the two slits? 0.442 mm 0o2 m C) 0.663 mm D) 0.221 mm E) 0.056 mm Problem 20: (5.00 pts) ie this as elnc intly oo am. Whath rf phodton A) 7.95 x 10-16J B) 1.60 x 10-15 J 3.98 x 1016J D) 3.18 x 10-15J E) 1.00 x 1016J
Answer #1

1] Let the speed of sound in the metal be V*L = 4 m

Time of travel is L/343 for air and L/V for metal

L(1/343 - 1/V) = 11*10^-3 s

(1/343 - 1/V ) = 11*10^-3/4

1/V = (1/343) - (11*10^-3/4)

V = 1/[(1/343) - (11*10^-3/4)] = 6044 m/s

No option is correct